# Enhancing the hyperpolarizability of crystals with quantum geometry

### v1

@@ Abstract
Discussion and conclusions.— In summary, we have shown that the colossal linear and higher-order optical responses observed in one-dimensional $\pi$-conjugate chains can be understood by means of the modern notions of band topology and quantum geometry, where the latter is bounded from below by the former. We have established a transparent connection of the multiband quantum geometry with the third-order optical responses, demonstrating that the crystalline-symmetry-protected topology induces a lower-bound on the hyperpolarizability. Our quantum-geometric perspective sheds a new light on the longstanding controversy of discrepancy betweeen $\chi^{(3)}$ of cis- and trans-polyacetylene  Bredas et al. (1994). In the provided picture, cis-polyacetylene is topologically trivial (${\nu_{\text{FK}}=0}$), hence, realizes no bound on geometry with average quantum metric $\langle g_{xx}\rangle$ that induces an enhancement of the $\chi^{(3)}$, unlike in the topological trans-polyacetylene ($\nu_{\text{FK}}=1$). To conclude, our work demonstrates that higher-order optical, or electrodynamical, responses can be achieved in materials possessing nontrivial crystalline topology, thus opening a pathway to identify and engineer novel platforms for enhanced nonlinear phenomena within the growing family of topologically classified materials.

### v2

@@ Abstract
Discussion and conclusions.— In summary, we have shown that the colossal linear and higher-order optical responses observed in one-dimensional $\pi$-conjugate chains can be understood by means of the modern notions of band topology and quantum geometry, where the latter is bounded from below by the former. We have established a transparent connection of the multiband quantum geometry with the third-order optical responses, demonstrating that the crystalline-symmetry-protected topology induces a lower bound on the hyperpolarizability. Our quantum-geometric perspective sheds a new light on the long-standing controversy of thediscrepancy betweeen $\chi^{(3)}$ of the cis and trans geometric isomers of polyacetylene  Sinclair et al. (1988); Yu et al. (1989b); Bredas et al. (1994). The distinction between cis- and trans-polyacetylene arises from the different spatial arrangements of atoms with identical chemical connectivity, which induces inequivalent band topology in the two isomers. The band topology classified by the winding of the Bloch eigenvectors provides for a lower bound in quantum geometric tensor and in the associated enhanced nonlinear optical response matrix elements in the trans isomer. While being manifestly larger in the topological polymer, the metric band average lies above the topological lower bound and does not saturate, unless the band dispersion is forced to approach the flat band limit. The minimization of lower-bounded quantum metric in such limiting case is consistent with the other known flatness conditions for the metric bound saturations Roy (2014); Kwon and Yang (2024); Wahl et al. (2025). In the provided picture, cis-polyacetylene is topologically trivial (${\nu_{\text{FK}}=0}$), hence, realizes no bound on geometry with average quantum metric $\langle g_{xx}\rangle$ that induces an enhancement of the $\chi^{(3)}$, unlike in the topological trans-polyacetylene ($\nu_{\text{FK}}=1$).
The theory presented in this work offers a step forward by revealing that the origin of the $p_{z}$-orbital band polarizability is topologically-induced, as associated with the lower-bounded quantum metric providing for an enhanced spread of the electronic wavefunction in the topological phase. To the best of our knowledge, the importance of band topology and geometry for nonlinear responses in the static limits— especially concerning hyperpolarizabilities—-were not addressed in previous studies. Given that any band topology is fully encoded in the non-Abelian Berry connections, which are fully captured by the quantum geometry determining nonlinear optical responses Ahn et al. (2021), our conceptual framework consistently and comprehensively captures the role of band topology for hyperpolarizabilities in any class of materials. In addition to offering an understanding of the observed difference in the hyperpolarizabilities between distinct isomers through the notion of quantum geometric bound, the phenomenological picture underlying our findings translates into topologically-induced quantum lattice fluctuations Yu et al. (1989a); Takahashi (1995, 1997b) in the form of virtual soliton-antisoliton pairs Yu et al. (1989a). Following identifications of positional charge correlators with electronic localization lengths and quantum geometry of Refs. Kivelson (1982); Souza et al. (2000); Tam et al. (2024), such quantum charge density fluctuations are driven by quantum metric.
To conclude, our work demonstrates that topologically lower-bounded higher-order geometric optical responses can be achieved in materials possessing nontrivial crystalline topology, thus opening a pathway to identify and engineer novel platforms for enhanced nonlinear phenomena within the growing family of topologically classified materials.